# On Selection of Cross-Section Averages in Non-stationary Environments

### v2

@@ 6.6.4 Correlated Idiosyncratics
Table 8: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$ with generated factors, $\overline{\mathbf{Z}}_{M}=\mathbf{F}\mathbf{q}_{M}$. DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2). For $DVS_{1}$ and $MW_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $DVS_{2}$ and $MW_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$. Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ are weakly correlated time and units units.
@@
Table 9: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$ with generated factors, $\overline{\mathbf{Z}}_{M}=\mathbf{F}\mathbf{q}_{M}$. DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2). For $DVS_{1}$ and $MW_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $DVS_{2}$ and $MW_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$. Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ are weakly correlated time and units units.
@@ 6.6.5 Eigenvalue Ratio
$\tau=0$ $\tau=0.1$ $\tau=0.9$
$N$ $T$ $ER(\mathbf{X})$ $ER(\widetilde{\mathbf{X}})$ $ER(\mathbf{Z})$ $ER(\widetilde{\mathbf{Z}})$ $ER(\mathbf{X})$ $ER(\widetilde{\mathbf{X}})$ $ER(\mathbf{Z})$ $ER(\widetilde{\mathbf{Z}})$ $ER(\mathbf{X})$ $ER(\widetilde{\mathbf{X}})$ $ER(\mathbf{Z})$ $ER(\widetilde{\mathbf{Z}})$
10 10 6.80 7.12 7.87 5.98 6.79 7.05 7.67 5.90 4.33 6.17 5.93 5.31
20 10 7.21 7.41 8.18 5.95 7.14 7.42 8.11 5.81 5.43 6.94 7.05 5.75
50 10 6.83 7.41 8.05 4.72 6.73 7.40 7.99 4.63 4.57 6.85 6.20 4.85
100 10 6.71 7.41 7.62 3.88 6.69 7.40 7.73 3.86 3.98 6.34 5.67 3.76
150 10 6.33 7.12 7.60 4.15 6.28 7.09 7.54 4.26 3.48 5.61 5.60 4.12
200 10 6.02 6.68 7.73 4.11 5.97 6.62 7.65 4.07 3.09 4.75 5.27 3.78
10 20 3.14 6.95 2.19 4.46 3.11 6.84 2.12 4.37 2.92 4.76 1.90 3.47
20 20 6.01 7.62 3.41 3.00 5.94 7.62 3.58 2.99 3.48 6.87 2.38 3.07
50 20 4.96 7.64 2.52 3.69 4.83 7.63 2.75 3.67 2.54 6.46 1.86 2.75
100 20 4.14 7.33 3.35 4.00 4.05 7.37 3.24 3.89 1.57 5.73 1.48 3.19
150 20 3.81 6.02 3.43 3.78 3.95 6.09 3.38 3.76 1.54 3.97 1.54 3.55
200 20 3.99 4.92 3.69 3.92 4.02 4.90 3.82 3.97 1.43 3.06 1.68 3.40
10 50 4.46 7.84 1.20 2.88 4.32 7.84 1.24 2.85 2.61 6.26 1.33 2.82
20 50 6.90 7.94 1.66 2.92 6.88 7.93 1.70 2.88 4.16 7.35 1.73 2.89
50 50 3.97 7.82 2.64 3.60 4.00 7.83 2.81 3.52 1.95 7.36 1.53 3.15
100 50 3.69 5.26 2.66 4.51 3.55 5.55 2.58 4.35 1.63 4.66 1.39 2.96
150 50 4.00 4.59 3.19 4.13 3.95 4.85 3.01 4.16 1.43 3.46 1.34 2.71
200 50 4.00 4.02 3.28 2.92 4.00 4.03 3.23 2.97 1.29 2.50 1.37 2.80
10 100 3.00 7.87 3.39 3.95 2.98 7.87 3.28 3.81 2.65 6.09 1.51 3.42
20 100 6.03 7.99 1.19 3.71 6.27 7.99 1.23 3.53 3.91 7.38 1.53 2.84
50 100 3.92 7.98 1.55 3.38 4.00 7.98 1.78 3.41 2.42 7.42 1.42 3.05
100 100 4.00 5.80 2.44 3.00 4.00 5.37 2.36 2.95 1.67 5.22 1.49 2.66
150 100 4.00 4.00 2.64 5.50 4.00 4.00 2.64 5.42 1.67 3.40 1.72 2.46
200 100 4.00 4.00 4.08 4.76 4.00 4.00 4.08 4.24 1.54 2.70 1.59 2.23
10 150 4.00 6.42 3.18 3.24 4.00 6.92 3.08 3.38 2.21 5.25 1.47 2.60
20 150 6.42 8.00 1.05 4.77 6.49 8.00 1.12 4.57 5.17 7.41 1.66 3.23
50 150 3.60 7.91 2.47 2.98 3.96 7.90 2.41 2.97 3.29 6.54 1.37 2.84
100 150 4.00 4.52 2.87 4.89 4.00 4.29 2.73 4.92 1.66 5.51 1.67 2.36
150 150 4.00 4.00 2.78 5.53 4.00 4.00 2.57 5.45 1.78 3.75 1.82 2.57
200 150 4.00 4.00 4.51 5.24 4.00 4.00 4.47 5.23 1.79 2.88 1.69 2.34
10 200 4.00 7.98 3.52 3.46 4.00 7.97 3.54 3.37 2.22 6.99 1.83 2.91
20 200 4.06 8.00 2.69 3.77 4.12 8.00 2.79 3.65 4.38 7.42 1.90 2.81
50 200 4.00 8.00 2.05 4.32 4.00 8.00 2.05 3.97 3.77 7.62 1.59 2.77
100 200 4.00 4.24 2.05 5.25 4.00 4.23 1.94 5.30 1.93 5.17 1.63 2.49
150 200 3.99 3.95 2.25 4.33 3.97 3.89 2.36 3.79 1.77 4.24 1.49 1.79
200 200 4.00 4.00 4.26 5.50 4.00 4.00 4.29 5.46 1.78 2.64 1.76 2.48

### v3

@@ 5.4.4 Correlated Idiosyncratics
Stationary $\mathbf{F}$ Non-Stationary $\mathbf{F}$
$\tau=0$ $\tau=0.4$ $\tau=0.9$
NN $T$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$
Correct Selection Frequency for $g$
50 50 74.10 28.50 98.70 72.70 11.70 6.40 12.30 4.30 0.60 0.10 0.00 0.00
100 50 94.80 93.30 99.40 97.10 94.00 89.00 99.30 97.20 23.50 15.90 43.20 26.90
200 50 96.80 95.20 99.90 99.50 64.50 61.40 97.90 95.30 1.60 1.30 24.30 19.20
300 50 100.00 100.00 100.00 100.00 90.00 88.30 91.00 89.60 10.30 8.80 12.30 10.20
500 50 100.00 100.00 100.00 100.00 99.80 99.80 99.70 99.40 47.10 45.10 50.10 47.10
50 100 77.30 86.90 98.80 96.10 69.10 58.00 99.90 98.70 7.70 5.80 33.00 21.60
100 100 98.90 94.90 99.60 96.80 69.20 56.10 78.20 59.70 3.00 1.20 4.00 1.10
200 100 99.70 100.00 100.00 100.00 99.90 99.70 99.90 99.90 34.70 27.80 52.90 40.20
300 100 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 54.60 48.80 77.30 70.90
500 100 100.00 100.00 100.00 100.00 99.90 99.90 99.20 98.90 30.40 27.40 23.50 20.50
50 200 99.30 99.00 100.00 100.00 94.50 94.10 95.70 89.80 9.80 8.50 8.70 5.60
100 200 95.90 96.40 100.00 100.00 92.30 88.30 99.80 99.70 5.40 4.60 21.40 13.70
200 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 45.40 36.40 76.80 63.60
300 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 27.60 22.60 94.70 88.30
500 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 66.00 59.80 95.40 91.50
50 300 99.10 99.60 100.00 100.00 85.40 82.60 100.00 99.90 6.70 5.60 18.70 15.70
100 300 100.00 100.00 100.00 100.00 95.50 98.50 100.00 100.00 19.10 17.50 36.50 29.90
200 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 56.80 50.80 63.10 55.10
300 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 90.50 84.10 98.30 95.10
500 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 85.00 79.90 98.00 96.70
50 500 100.00 100.00 100.00 100.00 85.10 83.10 100.00 100.00 2.80 2.30 12.30 11.30
100 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 36.80 34.40 60.60 55.60
200 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 83.10 80.80 98.60 97.80
300 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 88.60 84.70 99.10 98.10
500 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 93.60 88.70 99.90 99.70
Average of $g$
50 50 3.57 2.11 3.99 3.53 2.65 2.21 2.76 2.24 1.63 1.47 1.65 1.43
100 50 4.00 3.91 3.99 3.95 3.94 3.84 3.99 3.97 2.52 2.26 3.10 2.71
200 50 3.98 3.96 4.00 3.99 3.62 3.58 3.96 3.91 2.19 2.12 2.30 2.14
300 50 4.00 4.00 4.00 4.00 3.90 3.88 3.91 3.90 2.62 2.55 2.81 2.72
500 50 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.99 3.27 3.23 3.36 3.30
50 100 4.22 4.01 3.99 3.95 3.68 3.39 4.00 3.98 2.17 2.00 2.72 2.40
100 100 3.99 3.92 4.00 3.96 3.60 3.33 3.78 3.54 1.51 1.35 1.99 1.73
200 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 2.98 2.80 3.26 2.98
300 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.43 3.33 3.74 3.64
500 100 4.00 4.00 4.00 4.00 4.00 4.00 3.99 3.99 3.03 2.94 3.15 3.10
50 200 4.00 3.99 4.00 4.00 3.98 3.95 3.94 3.86 1.93 1.85 1.74 1.62
100 200 4.00 3.97 4.00 4.00 3.92 3.87 4.00 4.00 2.11 2.03 2.21 1.98
200 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.25 3.05 3.66 3.42
300 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.11 3.01 3.89 3.75
500 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.43 3.32 3.93 3.88
50 300 4.01 4.00 4.00 4.00 3.85 3.83 4.00 4.00 2.08 2.02 2.36 2.26
100 300 4.00 4.00 4.00 4.00 4.04 4.01 4.00 4.00 2.81 2.73 2.89 2.74
200 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.51 3.41 3.62 3.51
300 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.87 3.77 3.98 3.95
500 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.85 3.79 3.98 3.96
50 500 4.00 4.00 4.00 4.00 3.84 3.82 4.00 4.00 1.85 1.81 2.04 1.99
100 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.03 2.97 3.43 3.33
200 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.80 3.76 3.98 3.97
300 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.87 3.82 3.99 3.98
500 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.92 3.84 4.00 4.00
@@
Table 7: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$. DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2) in the main text. For $DVS_{1}$ and $MW_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $DVS_{2}$ and $MW_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$. Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ are weakly correlated time and units units.
Stationary $\mathbf{F}$ Non-Stationary $\mathbf{F}$
$\tau=0$ $\tau=0.4$ $\tau=0.9$
NN $T$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$ $MW_{1}$ $MW_{2}$ $DVS_{1}$ $DVS_{2}$
Correct Selection Frequency for $g$
50 50 100.00 99.80 97.20 61.60 36.40 26.90 13.90 4.60 0.50 0.10 0.00 0.00
100 50 100.00 99.80 99.30 96.40 99.80 99.70 99.10 96.30 63.70 53.20 38.00 23.50
200 50 100.00 100.00 99.80 99.40 100.00 100.00 98.50 95.80 56.70 50.60 24.50 18.70
300 50 100.00 100.00 100.00 100.00 97.20 96.70 91.80 89.70 29.60 26.10 12.50 10.70
500 50 100.00 100.00 100.00 100.00 100.00 100.00 99.60 99.60 77.20 74.90 50.40 46.50
50 100 100.00 100.00 97.10 90.80 100.00 99.90 99.70 97.10 50.10 38.70 19.60 11.20
100 100 100.00 100.00 99.50 96.50 93.10 85.10 72.10 54.20 4.60 2.40 0.70 0.30
200 100 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.90 89.00 84.60 49.20 36.50
300 100 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 93.80 91.10 77.30 70.00
500 100 100.00 100.00 100.00 100.00 100.00 100.00 99.20 98.90 48.90 44.30 23.30 20.40
50 200 100.00 100.00 100.00 99.90 100.00 100.00 94.00 87.50 57.70 52.20 2.40 1.80
100 200 100.00 100.00 100.00 100.00 100.00 100.00 99.80 99.80 53.50 45.20 12.40 7.70
200 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 97.40 93.40 72.50 56.60
300 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.40 98.70 93.10 87.20
500 200 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.70 99.70 94.90 91.00
50 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.60 54.70 50.90 7.60 6.20
100 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 84.30 80.70 26.10 19.80
200 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 92.20 89.30 61.30 52.60
300 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.70 97.20 93.60
500 300 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.80 97.90 96.30
50 500 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.90 14.70 13.60 4.00 3.50
100 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 93.90 93.20 47.60 43.90
200 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.90 98.20 96.60
300 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 98.50 97.00
500 500 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 99.90 99.50
Average of $g$
50 50 4.00 4.00 3.97 3.33 3.29 3.11 2.77 2.23 2.02 1.83 1.59 1.37
100 50 4.00 4.00 3.99 3.94 4.00 4.00 3.99 3.96 3.51 3.32 3.00 2.61
200 50 4.00 4.00 4.00 3.99 4.00 4.00 3.97 3.93 3.17 3.03 2.41 2.23
300 50 4.00 4.00 4.00 4.00 3.97 3.97 3.92 3.90 3.22 3.17 2.83 2.75
500 50 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.75 3.72 3.37 3.30
50 100 4.00 4.00 3.97 3.88 4.00 4.00 4.00 3.97 3.18 2.93 2.50 2.20
100 100 4.00 4.00 4.00 3.96 3.93 3.85 3.71 3.49 2.03 1.80 1.70 1.50
200 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.86 3.79 3.21 2.94
300 100 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.93 3.90 3.74 3.63
500 100 4.00 4.00 4.00 4.00 4.00 4.00 3.99 3.99 3.48 3.42 3.15 3.09
50 200 4.00 4.00 4.00 4.00 4.00 4.00 3.91 3.81 3.17 3.04 1.56 1.51
100 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.18 2.99 2.10 1.92
200 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.97 3.91 3.60 3.31
300 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.99 3.98 3.87 3.75
500 200 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.94 3.88
50 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.32 3.24 2.08 2.01
100 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.76 3.70 2.75 2.61
200 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.92 3.89 3.60 3.49
300 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.97 3.93
500 300 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.98 3.96
50 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 2.31 2.27 1.81 1.78
100 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.93 3.92 3.20 3.12
200 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.98 3.96
300 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.98 3.97
500 500 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 4.00 3.99
@@
Table 8: IC with Cross-Sectional Averages -Average of $g$ with $m=4$ and $K\in\{8,9\}$ with generated factors, $\overline{\mathbf{Z}}_{M}=\mathbf{F}\mathbf{q}_{M}$. DVS criteria from De Vos and Stauskas, (2024), MW from Margaritella and Westerlund, (2023), see (2.1) and (2.2) in the main text. For $DVS_{1}$ and $MW_{1}$, $p_{N,T}=\frac{N+T}{NT}\ln(\frac{NT}{N+T})$; for $DVS_{2}$ and $MW_{2}$, $p_{N,T}=\frac{N+T}{NT}\ln(C^{2}_{N,T})$ with $C_{N,T}=\min(\sqrt{N},\sqrt{T})$. Idiosyncratics in $\mathbf{x}_{i,t}$, $\mathbf{v}_{i,t}$, and $\varepsilon_{i,t}$ are weakly correlated time and units units.
@@ 5.4.5 Eigenvalue Ratio
$\tau=0$ $\tau=0.4$ $\tau=0.9$
[…]
Correct Selection Frequency for $g$
50 50 66.80 66.80 14.20 14.20 5.60 5.60 3.10 3.10 0.00 0.00 0.00 0.00
100 50 99.80 99.80 51.10 51.10 95.80 95.80 37.80 37.80 0.80 0.80 1.20 1.20
200 50 97.40 97.40 24.20 24.20 52.10 52.10 11.00 11.00 0.20 0.20 0.10 0.10
300 50 100.00 100.00 17.40 17.40 80.90 80.90 14.90 14.90 1.70 1.70 0.90 0.90
500 50 99.20 99.20 5.90 5.90 83.10 83.10 7.80 7.80 7.90 7.90 3.30 3.30
50 100 100.00 100.00 52.70 52.70 98.70 98.70 37.40 37.40 1.00 1.00 1.10 1.10
100 100 99.60 99.60 48.80 48.80 24.50 24.50 7.70 7.70 0.00 0.00 0.00 0.00
200 100 100.00 100.00 42.10 42.10 98.00 98.00 25.70 25.70 2.10 2.10 2.00 2.00
300 100 100.00 100.00 8.40 8.40 97.70 97.70 14.40 14.40 15.30 15.30 4.70 4.70
500 100 99.70 99.70 0.40 0.40 68.20 68.20 4.80 4.80 3.10 3.10 1.50 1.50
50 200 100.00 100.00 32.70 32.70 93.60 93.60 27.30 27.30 0.00 0.00 0.00 0.00
100 200 100.00 100.00 21.20 21.20 94.50 94.50 25.80 25.80 0.00 0.00 0.00 0.00
200 200 100.00 100.00 38.50 38.50 98.70 98.70 25.70 25.70 5.30 5.30 2.70 2.70
300 200 100.00 100.00 1.70 1.70 93.50 93.50 7.60 7.60 7.00 7.00 2.20 2.20
500 200 100.00 100.00 0.00 0.00 80.40 80.40 2.20 2.20 7.70 7.70 2.30 2.30
50 300 100.00 100.00 40.70 40.70 99.60 99.60 32.30 32.30 0.00 0.00 0.90 0.90
100 300 100.00 100.00 34.20 34.20 99.70 99.70 31.00 31.00 2.40 2.40 1.40 1.40
200 300 100.00 100.00 29.70 29.70 100.00 100.00 26.40 26.40 19.40 19.40 5.40 5.40
300 300 100.00 100.00 2.70 2.70 99.80 99.80 10.70 10.70 17.80 17.80 3.80 3.80
500 300 100.00 100.00 0.00 0.00 97.50 97.50 0.30 0.30 26.90 26.90 5.90 5.90
50 500 100.00 100.00 82.80 82.80 100.00 100.00 46.40 46.40 0.10 0.10 0.30 0.30
100 500 100.00 100.00 28.30 28.30 100.00 100.00 36.20 36.20 6.80 6.80 3.80 3.80
200 500 99.10 99.10 16.60 16.60 100.00 100.00 26.10 26.10 37.40 37.40 7.40 7.40
300 500 100.00 100.00 0.70 0.70 100.00 100.00 6.30 6.30 32.80 32.80 8.00 8.00
500 500 100.00 100.00 0.00 0.00 94.70 94.70 0.20 0.20 17.80 17.80 3.10 3.10
Average of $g$
50 50 3.78 3.78 1.97 1.97 3.02 3.02 2.05 2.05 2.64 2.64 1.45 1.45
100 50 4.00 4.00 2.98 2.98 3.94 3.94 2.92 2.92 1.49 1.49 1.65 1.65
200 50 3.97 3.97 3.20 3.20 3.40 3.40 2.49 2.49 1.49 1.49 1.48 1.48
300 50 4.00 4.00 4.73 4.73 3.92 3.92 4.23 4.23 1.50 1.50 1.58 1.58
500 50 4.01 4.01 4.94 4.94 4.16 4.16 4.89 4.89 1.72 1.72 1.74 1.74
50 100 4.00 4.00 2.69 2.69 3.99 3.99 2.71 2.71 1.98 1.98 1.66 1.66
100 100 4.00 4.00 3.08 3.08 2.30 2.30 2.16 2.16 2.05 2.05 1.38 1.38
200 100 4.00 4.00 3.06 3.06 3.98 3.98 3.22 3.22 1.50 1.50 1.54 1.54
300 100 4.00 4.00 4.89 4.89 4.02 4.02 4.73 4.73 1.90 1.90 1.81 1.81
500 100 4.00 4.00 5.00 5.00 4.21 4.21 4.91 4.91 1.59 1.59 1.54 1.54
50 200 4.00 4.00 1.98 1.98 3.94 3.94 2.26 2.26 3.40 3.40 1.54 1.54
100 200 4.00 4.00 1.64 1.64 3.91 3.91 2.21 2.21 2.04 2.04 1.45 1.45
200 200 4.00 4.00 3.44 3.44 4.00 4.00 3.81 3.81 1.54 1.54 1.66 1.66
300 200 4.00 4.00 4.98 4.98 4.03 4.03 4.90 4.90 1.97 1.97 1.72 1.72
500 200 4.00 4.00 5.00 5.00 4.18 4.18 4.97 4.97 1.77 1.77 1.70 1.70
50 300 4.00 4.00 2.25 2.25 3.99 3.99 2.40 2.40 2.56 2.56 1.62 1.62
100 300 4.00 4.00 2.03 2.03 4.00 4.00 2.23 2.23 1.91 1.91 1.75 1.75
200 300 4.00 4.00 3.12 3.12 4.00 4.00 3.56 3.56 2.03 2.03 1.78 1.78
300 300 4.00 4.00 4.97 4.97 4.00 4.00 4.88 4.88 2.18 2.18 1.96 1.96
500 300 4.00 4.00 5.00 5.00 4.03 4.03 5.00 5.00 3.00 3.00 2.93 2.93
50 500 4.00 4.00 3.48 3.48 4.00 4.00 2.90 2.90 2.90 2.90 1.60 1.60
100 500 4.00 4.00 1.85 1.85 4.00 4.00 2.12 2.12 2.00 2.00 1.81 1.81
200 500 3.99 3.99 4.25 4.25 4.00 4.00 4.36 4.36 2.59 2.59 2.03 2.03
300 500 4.00 4.00 4.99 4.99 4.00 4.00 4.94 4.94 2.68 2.68 2.40 2.40
500 500 4.00 4.00 5.00 5.00 4.05 4.05 5.00 5.00 2.42 2.42 2.25 2.25